OpenStudy (anonymous):

What quadrant are the equations in? A. 3x + 2y = -15 x - 4y = -20 B. 10x - y = -30 x - y = -2 C. 2x - 5y = 10 x + 2y = 10 D. 3x + y = 9 -7x - 7y = -14

OpenStudy (anonymous):

go there for some accuarate help

OpenStudy (anonymous):

I don't have a library card #. That's not going to help me.

OpenStudy (anonymous):

240100447204

OpenStudy (anonymous):

try that

OpenStudy (anonymous):

lemme know what happens

OpenStudy (anonymous):

It worked but I don't see it helping me too much.

OpenStudy (anonymous):

ok you shouldve pressed the math section

OpenStudy (anonymous):

I did.

OpenStudy (anonymous):

ok then you fill out the information on the bottom it shouldve let you type in your qestion

OpenStudy (anonymous):

Would this be algebra I or II?

OpenStudy (anonymous):

algebra 2

OpenStudy (anonymous):

What subtopic?

OpenStudy (anonymous):

what are your choices

OpenStudy (anonymous):

would it be graphing functions?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I just need to know how to graph them.

OpenStudy (anonymous):

girl go the bottom of tha page and get a tutor a live tutor will help you out

OpenStudy (anonymous):

I don't like that site.

OpenStudy (anonymous):

lol ok then well their helpful i jsut uesd them rite now and they gave me the rite answer to everyting

OpenStudy (anonymous):

Brizy, here's something else you can also try. Go to http://fooplot.com. It will graph the equations for you and you can see what quadrants they are in. You do have to solve for y first. So your first equation in A would be: 3x + 2y = -15 2y = -3x - 15 y = (-3/2)x - (15/2) Then you'd enter everything right of the "=" sign on that website. Hope that helps.

OpenStudy (anonymous):

now thats coool im gonna try that

OpenStudy (anonymous):

Yeah that seems complicated :-(

OpenStudy (anonymous):

my god girl you are hard to please

OpenStudy (anonymous):

Lol I've herd that before :-)

OpenStudy (anonymous):

Haha. It does take some work, but that process should at least make it easier.

OpenStudy (anonymous):

ya try it

OpenStudy (anonymous):

oh shadowfiend was here

OpenStudy (shadowfiend):

Still is.

OpenStudy (anonymous):

aaaaahhhhhh

OpenStudy (anonymous):

have i metntioned you look wonderful today:)

OpenStudy (shadowfiend):

BrizyBaby—the easiest way to answer these question is usually to graph them, honestly.

OpenStudy (anonymous):

I don't know how to graph the equations....that's what I need help with.

OpenStudy (shadowfiend):

Got it! Now we're at the heart of the problem :)

OpenStudy (shadowfiend):

Let's start with A then. We have two equations: 3x + 2y = -15 x - 4y = -20

OpenStudy (anonymous):

well i give up nice meeting you thoug Brizzy

OpenStudy (anonymous):

I need a easy way to know how to graph them.

OpenStudy (anonymous):

You too aminah.love18

OpenStudy (shadowfiend):

I find it easiest to graph equations when they're in y = mx + b (slope-intercept) form. So first things first, we try to solve the equations for y: \[\begin{align} 3x + 2y &= -15\\ 2y &= -3x - 15\\ y &= \frac{-3x - 15}{2} = -\frac{3}{2}x - \frac{15}{2} \end{align}\] And the other one: \[\begin{align} x - 4y &= -20\\ -4y &= -20 - x\\ y &= \frac{-20 - x}{4} = \frac{-20}{4} - \frac{1}{4}x\\ y &= -5 - \frac{1}{4}x = -\frac{1}{4}x - 5 \end{align}\]

OpenStudy (shadowfiend):

So, we now have: y = -3/2 x - 15/2 y = -1/4x - 5 The first thing to notice is that both of these have a negative slope. Do you know what a negative slope means on a graph?

OpenStudy (anonymous):

I don't get how you did the equations but yeah a negative slope goes down left to right.

OpenStudy (shadowfiend):

Whoops. Sorry then, let's go back to that :) Let's start with 3x + 2y = -15. We're trying to solve for y. That means we need y alone. Remember that an equation doesn't change if we subtract the same thing from both sides. So first things first, we move the 3x to the right: 3x + 2y = -15 3x - 3x + 2y = -15 - 3x Subtract 3x from both sides, 3x - 3x is 0: 0 + 2y = -15 - 3x 2y = -15 - 3x (We leave off the 0.) Then we need to get rid of the 2. Dividing the same thing from both sides also leaves the equation unchanged, so we'll divide by 2: y = (-15 - 3x) / 2 y = -15/2 - 3/2 x Does that make more sense?

OpenStudy (anonymous):

yeah. So the points are -7.5 and -1.5?

OpenStudy (shadowfiend):

Close! :) This means that the line is, like we said, going top left to bottom right. It means for every 3 units down that it goes, it goes two units to the right. And it means that it meets the y axis at -7.5. A quick, semi-accurate drawing: